OpenStudy (anonymous):

Algebra 2 help please??! Evaluate the area under g(x) = –x2 – 3x + 7 for –4 ≤ x ≤ 1. A.35.58 units^2 B.44.64 units^2 C.35.83 repeating units^2 D.none of these

OpenStudy (anonymous):

Do you know how to integrate between limits?

OpenStudy (anonymous):

so \[Area = \int\limits_{-4}^{1} (–x ^{2} – 3x + 7) dx\]

OpenStudy (anonymous):

no, honesty, I just don't get this

OpenStudy (anonymous):

that didn't help.. anyone else?

OpenStudy (anonymous):

Have you seen integration before at all? If you have I can talk you through it, but if not then I obviously can't

OpenStudy (anonymous):

ya, im looking at the lesson rite now.. im still confused?

OpenStudy (anonymous):

okay so i can explain integration to you... the whole purpose of integrating a function is to find the area underneath a curve. so you're given g(x) = –x^2 – 3x + 7 which looks a bit like this  so in order to find the area underneath the curve (ie the lined bit) we have to integrate the formula of the curve so if we have a function \[\int\limits_{?}^{?}x ^{n}dx = \frac{ x ^{n+1} }{ n+1 }\] so if we integrate g(x)=–x^2 – 3x + 7 we have to increase the powers by 1, and divide by the new power so \[\int\limits_{-1}^{4} (–x^2 – 3x + 7) dx = \frac{ -x ^{3} }{ 3 }-\frac{ 3x ^{2} }{ 2 }+7x\] then you have to substitute in your x=4 then also work out x=-1 and subtract one from the other